OpenStudy (moongazer):

who knows the the formula of graham's law of effusion here? when time or distance is given

OpenStudy (moongazer):

I'll just check if my formula is correct :)

OpenStudy (anonymous):

i think it is \[\Huge \frac{r_{1}}{r_{2}}=\frac{\sqrt{d_{2}}}{\sqrt{d_{1}}}\]

OpenStudy (anonymous):

@sami-21 I actually think its r(M)½ = constant? But I could be wrong

OpenStudy (moongazer):

molar mass and time are given

OpenStudy (anonymous):

@gjhfdfg you are correct .

OpenStudy (moongazer):

I think I know how to answer it already :) anyway, thanks for your answers

OpenStudy (anonymous):

in terms of molar masses \[\Large \frac{r_{1}}{r_{2}}=\frac{\sqrt{M_{2}}}{\sqrt{M_1}}\]